OpenStudy (anonymous):

Simplify the expression, 4/(3x)^-2

OpenStudy (anonymous):

well

OpenStudy (anonymous):

Hey

OpenStudy (anonymous):

hello!

OpenStudy (anonymous):

Guys some help please

OpenStudy (uri):

Your twin @FriedRice :o

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

seriously though someone answer my question

OpenStudy (anonymous):

\[\frac{ 4 }{ (3x)^-2 } = 4(9x^2) = 36x^2\]

OpenStudy (anonymous):

that's what I got